A rectangular lot is 80 meters wide and 130 meters long.Give the length and width of another rectangular lot that has the same perimeter but a smaller area.



Answer by checkley77(12844)

80 by 130
p=2*80+2*130
p=160+260
p=420 is the perimeter.
a=80*130=10,400 is the area.
------------
60 by 150 new dimentions.
p=2*60+2*150
p=120+300
p=420 same perimeter.
a=60*150
a=9,000 smaller area.